A population consist of the data 10,9,6,20,16 and 24.costruct a sampling distribution of size 3.


1
Expert's answer
2021-06-02T18:03:19-0400

N=6

n=3

Number of samples "= \frac{N!}{n!(N-n)!} = \frac{6!}{3!(6-3)!}= \frac{120}{6}=20"



A sampling distribution